Find the area enclosed between the curve y = x2 + 2, X-axis, x = 3 and x = 4


1
Expert's answer
2021-04-25T17:43:03-0400

The area region enclosed by the curvesΒ "y=x^2+2" , x-axis and x=4 is given by

"area= \u222b_3^4ydx=\u222b_3^4(x^2+2)dx"

"=[x^3\/3+2x]_3^4"

"=[4^3\/3+2(4)]-[3^3\/3+2(4)]"

"=88\/3-15"


=14.333 square units of lengthΒ